OpenStudy (anonymous):

Evaluate \[ \lim_{x \to (\pi/2)^-}\: cos(x)sec(5x) \]

OpenStudy (anonymous):

\[\lim_{x\to\pi/2^-}\frac{\cos x}{\cos5x}=\frac{0}{0}\] Apply L'Hopital's rule, if you know it.

OpenStudy (anonymous):

Does the - in the limit affect the answer? I get 1/5